# Phase Transitions in Nonreciprocal Driven-Dissipative Condensates

### v1

@@ Abstract
Introduction.—Driven-dissipative many-body systems have emerged as a versatile platform for exploring novel nonequilibrium phenomena that are inaccessible in equilibrium. These systems exhibit a rich variety of exotic phases of matter and phase transitions, as well as dynamic behaviors unique to nonequilibrium settings [1]. Much of this interest is motivated by the emergence of programmable quantum simulators in various platforms such as exciton-polariton systems in semiconductors [2], ultracold atomic systems [3], photonic platforms [4, 5], and superconducting circuits [6], where both the coherent interactions and dissipation can be engineered to a degree.
Recent years have seen a surge of interest in nonreciprocally interacting many-body systems, especially the nonequilibrium phases that emerge therein. Prototypical examples often consist of multi-species systems where the interactions between different species are asymmetric [7, 8, 9, 10, 11, 12, 13, 14]. Nonreciprocal interactions have also been shown to give rise to exotic nonequilibrium phases of matter, such as long-range order in two spatial dimensions [15, 16, 17], time-crystalline order [10, 11, 12, 8, 14], nonequilibrium boundary modes [18], and chaotic phases [19]. These phenomena have attracted significant attention in diverse fields such as active matter [20], biology [21], and neural networks [22], with early explorations of analogous quantum systems also beginning to emerge [23, 24, 25, 26, 27].
Another example of nonreciprocity arises when a system exhibits preferred spatial directionality. A paradigmatic example is the Hatano-Nelson model [28, 29], a one-dimensional system governed by a non-Hermitian Hamiltonian with asymmetric left and right hopping rates. This asymmetry leads to exponential localization of all eigenstates at one boundary under open boundary conditions (OBC), while under periodic boundary conditions (PBC), the eigenstates remain extended—a phenomenon known as the non-Hermitian skin effect (NHSE) [30, 31, 32]. While recent studies have explored nonlinear extensions of such models [33, 34, 35, 36, 37, 38, 39, 27], the impact on phase transitions and the specific role of boundaries, particularly beyond the framework of non-Hermitian Hamiltonians, remains largely underexplored.
@@
In this work, we investigate the effect of nonreciprocity and the NHSE in the presence of interactions on nonequilibrium phase transitions. We focus on a mean-field treatment of driven-dissipative condensates in one dimension, described by a Lindblad master equation, and find that the steady-state phase diagram is highly sensitive to boundary conditions. While periodic boundary conditions lead to a single phase characterized by a finite momentum and frequency condensate, open boundaries result in a rich variety of static and dynamical phases. This includes exotic phase transitions associated with particle-hole symmetry breaking and critical exceptional points (CEPs) [40, 41, 42, 24, 43, 44, 25, 12, 13, 9], as well as phases with distinct bulk and edge dynamics. Our work suggests that the NHSE affects physics in a rich dynamic manner due to the interplay between nonreciprocity and nonlinearity.
@@
where $\Delta=i(\kappa-2\gamma)$, $J_{+}=-(J+\gamma e^{i\theta})$, $J_{-}=-(J-\gamma e^{-i\theta})$. The interplay between the coherent dynamics ($J$) and the correlated loss ($\gamma$) generates asymmetric hoppings rates $J_{+}\neq J_{-}$, favoring left- (right-) propagating modes when $\theta=0$ ($\theta=\pi$). The NHSE is reflected in the spectrum of Eq. 1, illustrated in Fig. 1(b,c,d). The PBC spectrum is an ellipse in the complex plane, whereas for OBC the spectrum always falls on a line [see Fig. 1(b,c)]. For $\theta=\pm\pi/2$, the model is reciprocal (i.e. $|J_{+}|=|J_{-}|$), and the ellipse shrinks into a line signaling the absence of the NHSE [see Fig. 1(d)].
The quadratic system without two-particle losses can be dynamically unstable; with these interaction terms ($\Gamma\neq 0$) the system is always stable and becomes genuinely many-body. Here, we study the mean-field dynamics of the Lindbladian, described by the equation of motion (EOM)
@@
Periodic boundary conditions.—We begin with a study of the steady-state solutions of Eq. 2 with PBC, i.e., $\alpha_{N+1}=\alpha_{1}$. One finds that for large $N$, the vacuum state $\alpha_{j}=0$ is unstable for any amount of pumping $\kappa>0$. This can be seen from the spectrum of Eq. 1 which is diagonalized in momentum space $H_{\mathrm{HN}}=\sum_{q}(i\kappa-i\gamma_{q}+\omega_{q})\outerproduct{q}{q}$ | where
@@
are the decay rate and energy for momentum $q=2\pi m/N,\,m=0,\cdots,N-1$, For any choice of $\theta$, the mode at $q+\theta=-\pi/2$ (which is a dark mode of all $\hat{L}^{(2)}_{j}$ dissipators) experiences no dissipation and can be excited with infinitesimal $\kappa$ (for $N\rightarrow\infty$).
@@
The nonlinearity in Eq. 2 couples the plane wave solutions of the linear system, and one might expect that full behavior of the nonlinear equations would thus be more complex. This is not the case: the time-dependent traveling wave states (i.e. finite momentum condensates)
@@
are solutions of Eq. 2 with PBC whenever $\gamma_{q}<\kappa$. Not all of these plane-wave solutions are stable: a momentum-$q$ traveling wave can become unstable against creation of pairs at momenta $q\pm k$, for some $k$ that depends on $\kappa$ and $\gamma$ [see Supplemental Material (SM) for details [54]]. Figure 2 shows the resulting stability diagram for $\theta=\pi$ and $\gamma/J=0.5$. For $\kappa\rightarrow 0$, only the $q^{*}$ mode such that $\gamma_{q^{*}}\rightarrow 0$ is stable ($q^{*}=\pi/2$ for $\theta=\pi$), as expected. As $\kappa\rightarrow\infty$, on the other hand, all the modes in a range of $\pi$ centered about $q^{*}=-\pi/2-\theta$ are stable (corresponding to all the right-moving modes for $\theta=\pi$). Numerically, we find these simple traveling waves appear to be the only possible behavior at long times. For PBC and $\theta=\pi$, we thus find that the system always condenses into a single finite momentum mode (chosen by the initial conditions from a range of possible positive momenta), with a frequency $\omega_{q}$ that is non-zero (unless $q=\pi/2$). Other choices of $\theta$ lead to qualitatively similar results, with everything in Fig. 2 shifted to $q^{*}=-\pi/2-\theta$. Analogous traveling wave solutions also appear in the complex Landau-Ginzburg equation [55, 56], the continuum version of the reciprocal $\theta=-\pi/2$ limit of our model, where the stable possible condensates are centered about $q^{*}=0$.
@@
Open boundary conditions.—For PBC, the angle $\theta$ (and hence the nonreciprocity $|J_{+}|-|J_{-}|$) does not play a significant role. In contrast, for OBC, the value of $\theta$ is crucial to the physics. At $\theta=0,\pi$, the hopping is maximally nonreciprocal, and the full nonlinear system has an antilinear $\mathbb{Z}_{2}$ particle-hole $\mathcal{PH}$ symmetry. At the level of the EOM Eq. 2, this symmetry implies that if $\alpha_{j}(t)$ is a solution, then so is $e^{i\pi j}\alpha_{j}^{*}(t)$. We will demonstrate below that this symmetry can be spontaneously broken for OBC. At the level of the linear $\Gamma=0$ theory, this symmetry transforms the quadratic Hatano-Nelson Hamiltonian as $(\mathcal{PH})H_{\mathrm{HN}}(\mathcal{PH})^{-1}=-H_{\mathrm{HN}}$ [57]. In contrast, for more general $\theta\neq 0,\pi$, the $\mathcal{PH}$ symmetry is explicitly broken, making nontrivial steady states time-dependent. Here, we focus on the $\mathcal{PH}$-symmetric, maximally nonreciprocal case of $\theta=0,\pi$, choosing $\theta=\pi$ for concreteness. The corresponding phase diagram is shown in Fig. 3(a).
@@
The first intriguing feature of the OBC phase diagram is that the vacuum solution, $\alpha_{j}=0$, becomes stable within a certain range of parameters, shown in yellow in Fig. 3(a). This is in sharp contrast to the PBC case, where the vacuum is always unstable for finite pumping. The linear stability of the OBC vacuum phase is determined by examining the imaginary parts of the eigenvalues of Eq. 1 with OBC. The spectrum of Eq. 1 differs qualitatively depending on whether $\gamma<J$ or $\gamma>J$ [see Fig. 1(b, c)], with $\gamma=J$ marking an $N$-th order exceptional point. The eigenvalues are given by
@@
For $\gamma>J$, as $\kappa$ is increased above $\kappa_{\mathrm{crit}}$, the system undergoes a continuous phase transition from the vacuum phase ($\expectationvalue{\alpha_{j}}_{j}=0$) to a time-independent static condensate [blue region in Fig. 3(a)] with $\expectationvalue{\alpha_{j}}_{j}\neq 0$ [Fig. 3(c)], where $\expectationvalue{o}_{j}$ denotes the spatial average of $o$. Up to a spontaneously chosen global $U(1)$ phase, the complex amplitudes in the static phase are given by $\alpha_{j}^{s}=r_{j}^{s}e^{iq_{s}j}$, where $q_{s}=\pi/2$ ($-\pi/2$) for $\theta=\pi$ ($\theta=0$). Interestingly, this phase order corresponds to the minimal-damping wavevector of the PBC system in Eq. 3. This finite-momentum condensate emerges as $\kappa$ increases in the form of a kink that sweeps across the system from right to left [see Fig. 3(b)]. Numerically, we find that the height of the kink is $\sqrt{\kappa/\Gamma}$. In fact, the uniform state $\alpha_{j}=\sqrt{\kappa/\Gamma}e^{i\pi j/2}$ solves Eq. 2 (with $\theta=\pi$) for all $j$ except at the edges ($j=1,N$). As $\kappa\to\infty$, the system approaches this uniform state.
@@
Just above the phase transition, the kink first appears at the right edge ($j=N$) – consistent with the fact that for $\theta=\pi$, the linear eigenstates in the vacuum phase are all localized at the right boundary. As $\kappa$ increases further, the kink shifts toward the left boundary, leading to a continuous increase in the average amplitude $\expectationvalue{r_{j}}_{j}$ [Fig. 3(c)]. Numerically, we find that the kink position scales as $(\kappa-\kappa_{\mathrm{crit}})^{\beta}$ with $\beta=-0.5$ near the phase transition [see inset in Fig. 3(c)]. This critical exponent can be understood as follows: nonreciprocity induces directional propagation toward the right (for $\theta=\pi$). This net current balances with the local pump and loss, saturating the right-most sites at an amplitude of $\sqrt{\kappa/\Gamma}$. Disregarding these saturated sites, the remaining sites with zero amplitude form an effective smaller system. The finite-size expression in Eq. 5 can then be used to determine the new, slightly larger $\kappa$ required for the vacuum of this smaller system to become unstable. Alternatively, solving Eq. 5 for the effective system size – or equivalently the kink position – as a function of the distance from the phase transition ($\kappa-\kappa_{\mathrm{crit}}$) yields precisely $\beta=-0.5$.
@@
Next, we turn to the green region of the phase diagram in Fig. 3(a), which corresponds to time-dependent (dynamical) phases. This region has multiple attractors that encompasses a variety of phases and diverse dynamical behaviors. We will primarily focus on the most prominent of such phases, one that exists in almost the entire region. For fixed $\gamma/J<1$, dynamical phases exists only within a finite range of $\kappa$: they transition to the vacuum phase at small $\kappa$ and become unstable to the static condensate phase at large $\kappa$. To locate this upper phase boundary, we perform a numerical linear stability analysis of the static condensate state, yielding the red line in Fig. 3(a). Notably, the static condensate state $\alpha_{j}^{s}$ is $\mathcal{PH}$ symmetric (i.e., $\mathcal{PH}[\alpha_{j}^{s}]=\alpha_{j}^{s}$).
@@
Decreasing $\kappa$ along the leftmost white dashed line in Fig. 3(a) towards the red line, the dynamical matrix describing fluctuations of the static condensate develops a second zero eigenvalue, whose eigenvector then coalesces with the Goldstone mode to yield an instability. We thus identify this portion of the red line as a second order CEP (see SM [54]). Below the red line of Fig. 3(a), the system spontaneously breaks the $\mathcal{PH}$ symmetry, giving rise to two bistable time-dependent traveling wave states (wavevectors $q_{\pm}=q_{s}\pm\epsilon$) analogous to the PBC solutions in Eq. 4. The wavevectors $q_{\pm}$ evolves continuously as $\kappa$ is lowered, bifurcating from $q_{\pm}=q_{s}=\pi/2$ in the static phase (i.e. $\epsilon=0$) to either $q_{-}=0$ or $q_{+}=\pi$ at the point where the system enters the vacuum phase (i.e. $\epsilon=\pi/2$). Similarly, as shown in Fig. 3(d), the frequency evolves continuously from zero, bifurcating into two branches corresponding to clockwise (CW) or counterclockwise (CCW) rotations (analogous to the “chiral” states in Ref. [8]). The $\mathcal{PH}$ symmetry maps one solution onto the other, transforming a CW rotating state into a CCW one (i.e., flipping the sign of $\omega$), while simultaneously flipping the wavevector from $q_{s}+\epsilon$ to $q_{s}-\epsilon$, and vice versa. Remarkably, the PBC result for the frequency in Eq. 3 holds perfectly for these two states, using the numerically extracted average wavevector $\langle q\rangle_{j}$.
@@
Edge and bulk phase transitions.— So far, we have discussed cases where only the phase of each site was time-dependent. However, in some regions of parameter space within the dynamical phase, the amplitudes also becomes time-dependent. These amplitude fluctuations occur predominantly at the left edge of the system, resulting in distinct bulk versus edge behavior [see Fig. 4(a)]. In fact, the leftmost dashed-white line cut through the phase diagram of Fig. 3(a) already contains such an example. This is illustrated in Fig. 3(e), which displays the time- and space-averaged absolute value of the rate of change of the magnitude, $\langle|\partial_{t}r_{j}^{2}|\rangle_{j,t}$. This quantity is found to become nonzero at $\kappa/J\sim 1.9-2.2$. The inset in Fig. 3(e) shows an example of the time-averaged spatial profile ($\expectationvalue{|\partial_{t}r_{j}^{2}|}_{t}$) of these amplitude fluctuations, revealing that they are peaked at the left edge. Away from the edge, this quantity quickly drops, and then exponentially decays into the bulk. Notably, these amplitude fluctuations do not affect the $\langle\omega\rangle_{j}$ order parameter shown in Fig. 3(d), as the dominant Fourier component, even for the left-most sites, remains at the same bulk value as the average value $\langle\omega\rangle_{j}$ depicted in Fig. 3(d).
@@
The dynamical region ($\partial_{t}\alpha_{j}\neq 0$) in Fig. 4(a) includes additional phases with distinct edge and bulk behaviors, with some regions exhibiting multistability, as indicated by stripes in the phases in Fig. 4(a). Phase IV, indicated in purple in Fig. 4(a), features amplitude fluctuations at the left edge and phase-only oscillations in the bulk. A representative example (for $\gamma=0.3,\kappa=2.2$) of the amplitude and phase space-time dynamics is shown in Fig. 4(b,i) and Fig. 4(b,ii), respectively. Left of the black dashed line, the two limit cycles of phase I – the CW and the CCW solutions – merge, resulting in more complex behavior. The phases of the bulk sites act as pendula, involving both CW and CCW rotations [see Fig. 4(e)]. In some parts of this phase, particularly at smaller $\gamma$, the dynamics is nevertheless periodic. Notably, the $\mathbb{Z}_{2}$ symmetry is dynamically restored in this case, corresponding to $\mathcal{PH}$ accompanied by the breaking of continuous time translation symmetry [58]. Specifically, the $\mathcal{PH}$ operation acts as a translation by a half period, and the steady-state satisfies $\alpha_{j}(t)=\mathcal{PH}[\alpha_{j}(t+T/2)]$ for some period $T$ (see SM [54] for details).
@@
In Phase IV, the steady state alternates between exhibiting periodic and chaotic dynamics until a critical value of $\gamma$ where only the chaotic dynamics persists. The transition occurs in the purple region to the right of the black dashed line in Fig. 4(a). Surprisingly, we find that despite the onset of chaos, the dynamics at the edge remain regular (see SM [54] for additional details). We also find a region [Phase III in Fig. 4(a)] with a hierarchy of steady states with varying edge lengths. A representative example ($\gamma=0.4,\kappa=1.25$) is shown in Fig. 4(c). The bulk dynamics is $\mathcal{PH}$ broken, with CW or CCW phase oscillations [Fig. 4(e)] as in Phase I. We observe five steady states with progressively longer edge-amplitude fluctuations, shown in Fig. 4(d), although it remains unclear whether additional steady states exist. While the bulk dynamics in Phases II and III are periodic, the edge exhibits quasiperiodicity, breaking time translational symmetry and introducing an additional zero mode (beyond the existing Goldstone mode from a broken $U(1)$ symmetry, see SM [54]). Exploring these edge phase transitions in greater detail is left for future work.
@@
Conclusions and outlook.—We have demonstrated how the NHSE and spatial nonreciprocity can fundamentally alter the phase diagram of driven-dissipative systems in a manner that strongly depends on boundary conditions. The remarkable differences between OBC and PBC seen here are in stark contrast to the behavior of the fermionic and spin counterparts of our model [48, 27]. In those systems, excitations generically have a finite lifetime; they thus damp out before hitting the edge, making bulk properties insensitive to boundaries. In contrast, our bosonic system can possess an undamped zero mode (i.e. the Goldstone mode associated with the spontaneous $U(1)$ symmetry breaking). Such excitations can propagate to the boundaries, allowing a strong boundary-condition dependence.

### v2

@@ Abstract
Introduction.—Driven-dissipative many-body systems have emerged as a versatile platform for exploring novel nonequilibrium phenomena that are inaccessible in equilibrium. These systems exhibit a rich variety of exotic phases of matter, phase transitions, and unique dynamical behaviors [1]. Much of the interest is driven by the rise of programmable quantum simulators across various platforms, including exciton-polariton systems in semiconductors [2], ultracold atomic systems [3], photonic platforms [4, 5], and superconducting circuits [6], where both coherent interactions and dissipation can be engineered.
Recent years have seen a surge of interest in nonreciprocally interacting many-body systems. Prototypical examples often consist of multi-species systems where the interactions between different species are asymmetric [7, 8, 9, 10, 11, 12, 13, 14]. Nonreciprocal interactions have also been shown to give rise to exotic nonequilibrium phases of matter, such as long-range order in two spatial dimensions [15, 16, 17], time-crystalline order [10, 11, 12, 8, 14], nonequilibrium boundary modes [18], and chaotic phases [19]. These phenomena have attracted significant attention in diverse fields such as active matter [20], biology [21], and neural networks [22], with explorations of analogous quantum systems also beginning to emerge [23, 24, 25, 26, 27].
Another example of nonreciprocity arises when a system exhibits preferred spatial directionality [28, 29, 30, 31, 32, 33]. A paradigmatic example is the Hatano-Nelson model [34, 35], a one-dimensional system governed by a non-Hermitian Hamiltonian with asymmetric left and right hopping rates. This asymmetry causes exponential localization of all eigenstates at one boundary under open boundary conditions (OBC), while under periodic boundary conditions (PBC), the eigenstates remain extended—a phenomenon known as the non-Hermitian skin effect (NHSE) [36, 37, 38]. While recent studies have explored nonlinear extensions of such models [39, 40, 41, 42, 43, 44, 31, 27], the impact on phase transitions and the specific role of boundaries, particularly beyond the framework of non-Hermitian Hamiltonians, remains largely underexplored.
@@
In this work, we investigate the effect of nonreciprocity and the NHSE in the presence of interactions on nonequilibrium phase transitions. We focus on a mean-field treatment of driven-dissipative condensates in one dimension, described by a Lindblad master equation, and find that the steady-state phase diagram is highly sensitive to boundary conditions. While periodic boundary conditions lead to a single phase characterized by a finite momentum and frequency condensate, open boundaries result in a rich variety of static and dynamical phases. This includes exotic phase transitions associated with particle-hole symmetry breaking and critical exceptional points (CEPs) [45, 46, 47, 24, 48, 49, 25, 12, 13, 9], as well as phases with distinct bulk and edge dynamics.
@@
where $\Delta=i(\kappa-2\gamma)$, $J_{+}=-(J+\gamma e^{i\theta})$, $J_{-}=-(J-\gamma e^{-i\theta})$. The interplay between the coherent dynamics ($J$) and the correlated loss ($\gamma$) generates asymmetric hoppings rates $J_{+}\neq J_{-}$, favoring left- (right-) propagating modes when $\theta=0$ ($\theta=\pi$). The NHSE is reflected in the spectrum of Eq. 1, illustrated in Fig. 1(b,c,d). The PBC spectrum is an ellipse in the complex plane, whereas for OBC the spectrum always falls on a line [Fig. 1(b,c)]. For $\theta=\pm\pi/2$, the model is reciprocal (i.e., $|J_{+}|=|J_{-}|$), and the ellipse shrinks into a line signaling the absence of the NHSE [Fig. 1(d)].
The quadratic system ($\Gamma=0$) can be dynamically unstable; with these interaction terms ($\Gamma\neq 0$) the system is always stable and becomes genuinely many-body. Here, we study the mean-field dynamics of the Lindbladian, described by the equation of motion (EOM)
@@
Periodic boundary conditions.—We begin with a study of the steady-state solutions of Eq. 2 with PBC, i.e., $\alpha_{N+1}=\alpha_{1}$. For large $N$, the vacuum state $\alpha_{j}=0$ is unstable for any $\kappa>0$. This can be seen from the spectrum of Eq. 1 which is diagonalized in momentum space $H_{\mathrm{HN}}=\sum_{q}(i\kappa-i\gamma_{q}+\omega_{q})\outerproduct{q}{q}$ | where
@@
are the decay rate and energy for momentum $q=2\pi m/N,\,m=0,\cdots,N-1$. For any $\theta$, the mode at $q+\theta=-\pi/2$ (a dark mode of all $\hat{L}^{(2)}_{j}$ dissipators) experiences no dissipation and can be excited with infinitesimal $\kappa$ (for $N\rightarrow\infty$).
@@
Despite the nonlinearity in Eq. 2 coupling these plane waves, the time-dependent traveling wave states (i.e. finite momentum condensates)
@@
are solutions of Eq. 2 with PBC whenever $\gamma_{q}<\kappa$. Not all of these plane-wave solutions are stable: a momentum-$q$ traveling wave can become unstable against creation of pairs at momenta $q\pm k$, for some $k$ that depends on $\kappa$ and $\gamma$ [see Supplemental Material (SM) for details [59]]. Figure 2 shows the resulting stability diagram for $\theta=\pi$ and $\gamma/J=0.5$. For $\kappa\rightarrow 0$, only the $q^{*}$ mode such that $\gamma_{q^{*}}\rightarrow 0$ is stable ($q^{*}=\pi/2$ for $\theta=\pi$), as expected. As $\kappa\rightarrow\infty$, on the other hand, all the modes in a range of $\pi$ centered about $q^{*}=-\pi/2-\theta$ are stable (corresponding to all the right-moving modes for $\theta=\pi$). Numerically, these traveling waves appear to be the only possible behavior at long times. For PBC and $\theta=\pi$, we thus find that the system always condenses into a single positive, finite momentum mode (chosen by the initial conditions), with a nonzero frequency $\omega_{q}$ (unless $q=\pi/2$). Other choices of $\theta$ lead to qualitatively similar results, with everything in Fig. 2 shifted to $q^{*}=-\pi/2-\theta$. Analogous traveling wave solutions also appear in the complex Landau-Ginzburg equation [60, 61], the continuum version of the reciprocal $\theta=-\pi/2$ limit of our model, where the stable possible condensates are centered about $q^{*}=0$.
@@
Open boundary conditions.—For PBC, the angle $\theta$ (and hence the nonreciprocity $|J_{+}|-|J_{-}|$) does not play a significant role. In contrast, for OBC, the value of $\theta$ is crucial to the physics. At $\theta=0,\pi$, the hopping is maximally nonreciprocal, and the full nonlinear system has an antilinear $\mathbb{Z}_{2}$ particle-hole $\mathcal{PH}$ symmetry. At the level of the EOM, Eq. 2, this symmetry implies that if $\alpha_{j}(t)$ is a solution, then so is $\mathcal{PH}[\alpha_{j}(t)]=e^{i\pi j}\alpha_{j}^{*}(t)$. We will demonstrate below that this symmetry can be spontaneously broken for OBC. At the level of the linear $\Gamma=0$ theory, this symmetry transforms the quadratic Hatano-Nelson Hamiltonian as $(\mathcal{PH})H_{\mathrm{HN}}(\mathcal{PH})^{-1}=-H_{\mathrm{HN}}$ [62]. In contrast, for more general $\theta\neq 0,\pi$, the $\mathcal{PH}$ symmetry is explicitly broken, making nontrivial steady states time-dependent. Here, we focus on the $\mathcal{PH}$-symmetric, maximally nonreciprocal case of $\theta=0,\pi$, choosing $\theta=\pi$ for concreteness. The corresponding phase diagram is shown in Fig. 3(a).
@@
The first intriguing feature of the OBC phase diagram is that the vacuum solution, $\alpha_{j}=0$, becomes stable within a certain range of parameters, shown in yellow in Fig. 3(a). This is in sharp contrast to the PBC case, where the vacuum is always unstable for $\kappa>0$. The linear stability of the OBC vacuum phase is determined by examining the imaginary parts of the eigenvalues of Eq. 1 with OBC. The spectrum of Eq. 1 differs qualitatively depending on whether $\gamma<J$ or $\gamma>J$ [Fig. 1(b, c)], with $\gamma=J$ marking an $N$-th order exceptional point. The eigenvalues are given by [59]
@@
For $\gamma>J$, as $\kappa$ is increased above $\kappa_{\mathrm{crit}}$, the system undergoes a continuous phase transition from the vacuum phase ($\expectationvalue{\alpha_{j}}_{j}=0$) to a time-independent static condensate [blue region in Fig. 3(a)] with $\expectationvalue{\alpha_{j}}_{j}\neq 0$ [Fig. 3(c)], where $\expectationvalue{o}_{j}$ denotes the spatial average of $o$. Up to a spontaneously chosen global $U(1)$ phase, the complex amplitudes in the static phase are given by $\alpha_{j}^{s}=r_{j}^{s}e^{iq_{s}j}$, where $q_{s}=\pi/2$ ($-\pi/2$) for $\theta=\pi$ ($\theta=0$). Interestingly, this phase order corresponds to the minimal-damping wavevector of the PBC system in Eq. 3. This finite-momentum condensate emerges as $\kappa$ increases in the form of a kink that sweeps across the system from right to left [Fig. 3(b)]. Numerically, we find that the height of the kink is $\sqrt{\kappa/\Gamma}$. In fact, the uniform state $\alpha_{j}=\sqrt{\kappa/\Gamma}e^{i\pi j/2}$ solves Eq. 2 (with $\theta=\pi$) for all $j$ except at the edges ($j=1,N$). As $\kappa\to\infty$, the system approaches this uniform state.
@@
Just above the phase transition, the kink first emerges at the right edge ($j=N$), reflecting the localization of linear eigenstates at this boundary for $\theta=\pi$ in the vacuum phase. As $\kappa$ increases further, the kink shifts toward the left boundary, leading to a continuous increase in the average amplitude $\expectationvalue{r_{j}}_{j}$ [Fig. 3(c)]. Numerically, we find that the kink position scales as $(\kappa-\kappa_{\mathrm{crit}})^{\beta}$ with $\beta=-0.5$ near the phase transition [inset in Fig. 3(c)]. This critical exponent can be understood as follows: nonreciprocity induces directional propagation toward the right (for $\theta=\pi$). This net current balances with the local pump and loss, saturating the right-most sites at an amplitude of $\sqrt{\kappa/\Gamma}$. Disregarding these saturated sites, the remaining sites with zero amplitude form an effective smaller system. The finite-size expression in Eq. 5 can then be used to determine the new, slightly larger $\kappa$ required for the vacuum of this smaller system to become unstable. Alternatively, solving Eq. 5 for the effective system size – or equivalently the kink position – as a function of the distance from the phase transition ($\kappa-\kappa_{\mathrm{crit}}$) yields precisely $\beta=-0.5$.
@@
Next, we examine the green region of the phase diagram in Fig. 3(a), which corresponds to time-dependent (dynamical) phases. This region contains multiple attractors that encompass a variety of phases. We will primarily focus on the most prominent phase, which spans nearly the entire region. For fixed $\gamma/J<1$, dynamical phases exist only within a finite range of $\kappa$: they transition to the vacuum phase at small $\kappa$ and become unstable to the static condensate phase at large $\kappa$. To locate this upper phase boundary, we perform a numerical linear stability analysis of the static condensate state, yielding the red line in Fig. 3(a). Notably, the static condensate state $\alpha_{j}^{s}$ is $\mathcal{PH}$ symmetric (i.e., $\mathcal{PH}[\alpha_{j}^{s}]=\alpha_{j}^{s}$).
@@
Decreasing $\kappa$ along the leftmost white dashed line in Fig. 3(a) towards the red line, the dynamical matrix describing fluctuations of the static condensate develops a second zero eigenvalue, whose eigenvector coalesces with the Goldstone mode (associated with the spontaneous $U(1)$ symmetry breaking). We thus identify this portion of the red line as a second-order CEP [59]. Below the red line of Fig. 3(a), the system spontaneously breaks the $\mathcal{PH}$ symmetry, giving rise to two bistable time-dependent traveling wave states (wavevectors $q_{\pm}=q_{s}\pm\epsilon$) analogous to the PBC solutions in Eq. 4. The wavevectors $q_{\pm}$ evolves continuously as $\kappa$ is lowered, bifurcating from $q_{\pm}=q_{s}=\pi/2$ in the static phase ($\epsilon=0$) to either $q_{-}=0$ or $q_{+}=\pi$ at the point where the system enters the vacuum phase ($\epsilon=\pi/2$). Similarly, as shown in Fig. 3(d), the frequency evolves continuously from zero, bifurcating into two branches corresponding to clockwise (CW) or counterclockwise (CCW) rotations (analogous to the “chiral” states in Ref. [8]). The $\mathcal{PH}$ symmetry maps one solution onto the other, transforming a CW rotating state into a CCW one (i.e., flipping the sign of $\omega$), while simultaneously flipping the wavevector from $q_{s}+\epsilon$ to $q_{s}-\epsilon$, and vice versa. Remarkably, the PBC result for the frequency in Eq. 3 holds perfectly for these two states, using the numerically extracted average wavevector $\langle q\rangle_{j}$.
@@
Edge and bulk phase transitions.— So far, we have considered cases where only the phase of each site varied in time. However, in certain regions of the dynamical phase, the amplitudes also fluctuate. These fluctuations occur predominantly at the left edge, leading to distinct bulk versus edge behavior [Fig. 4(a)]. In fact, Fig. 3(d) already contains such an example. This is illustrated in Fig. 3(e), which displays the time- and space-averaged absolute value of the rate of change of the magnitude, $\langle|\partial_{t}r_{j}^{2}|\rangle_{j,t}$. This quantity becomes nonzero at $\kappa/J\sim 1.9-2.2$. The inset in Fig. 3(e) shows an example of the time-averaged spatial profile ($\expectationvalue{|\partial_{t}r_{j}^{2}|}_{t}$) of these amplitude fluctuations, revealing that they are peaked at the left edge. Away from the edge, this quantity quickly drops, and then exponentially decays into the bulk. Notably, these amplitude fluctuations do not affect the $\langle\omega\rangle_{j}$ order parameter shown in Fig. 3(d), as the dominant Fourier component, even for the left-most sites, remains at the same bulk value as the average value $\langle\omega\rangle_{j}$ depicted in Fig. 3(d).
@@
The dynamical region ($\partial_{t}\alpha_{j}\neq 0$) includes additional phases with distinct edge and bulk behaviors, with some regions exhibiting multistability, indicated by stripes in Fig. 4(a). Phase IV, indicated in purple in Fig. 4(a), features amplitude fluctuations at the left edge and phase-only oscillations in the bulk. A representative example ($\gamma=0.3,\kappa=2.2$) of the amplitude and phase space-time dynamics is shown in Fig. 4(b,i) and Fig. 4(b,ii), respectively. Left of the black dashed line in Fig. 4(a), the two limit cycles of Phase I – the CW and the CCW solutions – merge, resulting in more complex dynamics. The phases of the bulk sites act as pendula, involving both CW and CCW rotations [see Fig. 4(e)]. In some parts of this phase, particularly at smaller $\gamma$, the dynamics is nevertheless periodic. Notably, the $\mathbb{Z}_{2}$ symmetry is dynamically restored in this case, corresponding to $\mathcal{PH}$ accompanied by the breaking of continuous time translation symmetry [63, 64]. Specifically, the $\mathcal{PH}$ operation acts as a translation by a half period, and the steady state satisfies $\alpha_{j}(t)=\mathcal{PH}[\alpha_{j}(t+T/2)]$ for some period $T$ [59].
@@
In Phase IV, the steady state alternates between exhibiting periodic and chaotic dynamics until a critical value of $\gamma$ [black dashed line in Fig. 4(a)] beyond which only the chaotic dynamics persists. Surprisingly, we find that despite the onset of chaos, the dynamics at the edge remain regular [59]. We also find a region [Phase III in Fig. 4(a)] with a hierarchy of steady states with varying edge lengths. A representative example ($\gamma=0.4,\kappa=1.25$) is shown in Fig. 4(c). The bulk dynamics is $\mathcal{PH}$ broken, with CW or CCW phase oscillations [Fig. 4(e)] as in Phase I. We observe five steady states with progressively longer edge-amplitude fluctuations, shown in Fig. 4(d), although it remains unclear whether additional steady states exist. While the bulk dynamics in Phases II and III are periodic, the edge exhibits quasiperiodicity, breaking time translational symmetry and introducing an additional zero mode (beyond the existing Goldstone mode [59]). Exploring these edge phase transitions in greater detail is left for future work.
@@
Conclusions and outlook.—We have demonstrated how the NHSE and spatial nonreciprocity can fundamentally alter the phase diagram of driven-dissipative systems in a manner that strongly depends on boundary conditions. The remarkable differences between OBC and PBC seen here are in stark contrast to the fermionic and spin counterparts of our model [53, 27], where excitations generically have finite lifetimes and decay before reaching the edge, leaving bulk properties unchanged. In contrast, our bosonic system can possess an undamped Goldstone mode, allowing excitations to propagate to the boundaries and making boundary effects significant.